How many years will it take $100 to triple to $3.00 at an annual rate of 596 & Select one O a 28.5 O 0.615 O c.225 TO
Answer #1

Answer : (c) 22.50

Explanation

P ( 1 + i)n  = FV

Where P = Principal

i = Interest rate

FV = Future value

n= number of year

so, P ( 1 + i)n  = FV

or   $ 1 ( 1 + 0.05 ) n = $ 3

or ( 1.05 ) n = $ 3

n = 22.50 Years ( Approx )